### v1

Abstract
Surprising videos, e.g., funny clips, creative performances, or visual illusions, attract significant attention. Enjoyment of these videos is not simply a response to visual stimuli; rather, it hinges on the human capacity to understand (and appreciate) commonsense violations depicted in these videos. We introduce FunQA, a challenging video question answering (QA) dataset specifically designed to evaluate and enhance the depth of video reasoning based on counter-intuitive and fun videos. Unlike most video QA benchmarks which focus on less surprising contexts, e.g., cooking or instructional videos, FunQA covers three previously unexplored types of surprising videos: 1) HumorQA, 2) CreativeQA, and 3) MagicQA. For each subset, we establish rigorous QA tasks designed to assess the model’s capability in counter-intuitive timestamp localization, detailed video description, and reasoning around counter-intuitiveness. We also pose higher-level tasks, such as attributing a fitting and vivid title to the video, and scoring the video creativity. In total, the FunQA benchmark consists of 312K free-text QA pairs derived from 4.3K video clips, spanning a total of 24 video hours. Extensive experiments with existing VideoQA models reveal significant performance gaps for the FunQA videos across spatial-temporal reasoning, visual-centered reasoning, and free-text generation.
Figure 1: Overview of FunQA. FunQA comprises three subsets of surprising videos: 1) HumorQA, 2) CreativeQA, and 3) MagicQA. Each subset is associated with three common tasks: 1) counter-intuitive timestamp localization, 2) detailed video description, and 3) reasoning around counter-intuitiveness (see H1-3, C1-3, and M1-3). Furthermore, we offer higher-level tasks tailored for each video type, such as attributing a fitting and vivid title for HumorQA and CreativeQA (see H4, C4), etc.
1 Introduction
The charm of surprising videos, be they funny, creative, or filled with visual illusions, offer enjoyment and commands attention from viewers. This type of media elicits positive surprise [1],111c.f., negative surprise, e.g., a surprising medical bill. a captivating emotion that stems not merely from perceiving surface-level visual stimuli, but rather, the innate ability of humans to understand and find delight in unexpected and counter-intuitive moments [2]. However, despite significant advancements in today’s computer vision models, the question remains: can video models “understand" the humor/creativity in surprising videos?
Consider the humorous video depicted in Figure 1 (left) as an illustrative example. We witness a man engrossed in his phone, sharing a meal with friends. Suddenly, one of his companions squeezes a generous amount of ketchup, which, instead of adorning the fries, splatters onto the man’s face. The shock in his eyes, combined with his ketchup-covered visage, elicits laughter. [^2] The hostility/superiority theory of humor posits that humor can arise from claiming superiority over someone or something [3, 4]; but alternate (more optimistic) theories of humor exist, [5] offers a survey. While humans effortlessly recognize this as an unusual (and potentially entertaining) event, the reasoning required to holistically understand the scene is complex: a model needs to recognize that individuals were gathered to enjoy a meal together, and discern that the comedic element arises from the ketchup intended for the fries ending up on the man’s face instead, and that the innocent expression of the unfortunate victim indicates no significant harm was caused.
While there have been some efforts to enhance computer vision models’ performance in Video Question Answering (VideoQA), these works have primarily focused on the common, less surprising videos found in existing VideoQA datasets. Examples of commonly employed VideoQA datasets include YouCook2 [6] which contains video clips from 2K cooking videos, Howto100M [7] which consists of only instructional videos. While there exist video datasets that explore the humor in TV shows [8, 9] and include tasks such as predicting laughter tracks [10], these tasks often heavily rely on audio and narrative cues, with visual clues might playing a lesser role.
To address this gap and evaluate computer vision models’ ability to identify and understand visual commonsense violations in videos, we introduce FunQA, a comprehensive and high-quality VideoQA dataset comprising 4.3K surprising videos and 312K manually annotated free-text QA pairs. Our dataset consists of three subsets: 1) HumorQA, 2) CreativeQA, and 3) MagicQA. Each subset covers different sources and video contents, but the commonality lies in their surprising nature, e.g., the unexpected contrasts in humorous videos, the intriguing disguises in creative videos, and the seemingly impossible performances in magic videos. Our experiments suggest that these surprising videos require different types of reasoning than common videos, as existing VideoQA methods perform poorly on the corpus. With FunQA, we hope to provide a benchmark that covers the popular, important, and sophisticated genre of counter-intuitive/surprising videos.
In FunQA, we formulate three rigorous tasks to measure models’ understanding of surprise: 1) Counter-intuitive timestamp localization: a model must identify the specific time period within a video when an unexpected event takes place. 2) Detailed video description: a model must generate coherent and objective descriptions of the video content, evaluating models’ fundamental video understanding capabilities. 3) Counter-intuitiveness reasoning: a model must generate concrete explanations of why the video is surprising. These tasks progressively assess the model’s ability to perceive, articulate, and reason about the counter-intuitive elements present in surprising videos. Additionally, we propose auxiliary tasks that pose higher-level challenges within the benchmark including assigning an appropriate and vivid title to the video, etc. To summarize our contributions:
1) New VideoQA Dataset: We build a large-scale dataset FunQA, which complements the existing VideoQA dataset with intriguing videos.
2) Novel and Challenging Tasks: We design a number of novel tasks that allow the model to explore previously untouched problems, such as timestamp localization, and reasoning around counter-intuitiveness. These tasks push video reasoning beyond superficial descriptions, demanding deeper understanding and discernment.
3) Comprehensive Evaluation: We have done an extensive and comprehensive evaluation of cutting-edge baselines, giving the field an insight and future research direction.
2 Related Work
Table 1: Comparison between FunQA and other existing benchmarks. Compare to other datasets, FunQA revolves around the captivating realm of interesting and counter-intuitive videos. The tasks within FunQA are specifically designed to challenge the vision capabilities of models, requiring strong skills in producing an in-depth description, interpretation, and spatial-temporal reasoning. Here we clarify the abbreviation in the table. Anno.: Annotation; M: Manual, A: Automatic; For Input, V, A, S, and B denote Video, Audio, Subtitle, and Bounding-box; VC means visual-centric, Desc. means Description, Expl for Explanation, STR for Spatial-temporal Reasoning. For QA Tasks, MC denotes Multiple Choice QA, OE means Open Ended QA, and FT means Free Text QA.
Video Question Answer
Datasets Domain Anno. Avg length (s) # Clips (K) Input # QA pairs (K) VC Desc. Expl. STR QA Task
MarioQA [11] Games M 3-6 188 V, A 188 Yes No Yes Yes OE
TGIF-QA [12] Social Media M 3.1 71.7 V 165.2 Yes Yes No Yes MC & OE
MovieQA [13] Movies A 202.7 6.77 V, A, S 6.4 No No Yes Yes MC
CLEVRER [14] Synthetic Video M 5 20 V 305 Yes No Yes Yes OE
TVQA [15] TV shows M 76.2 21.8 V, A 152.5 No No Yes No MC
TVQA+ [16] TV shows M 7.2 4.2 V, A, B 29.4 No No Yes Yes OE
Social-IQ [17] Web videos M 99 1.25 V, A, S, B 7.5 Yes No Yes No MC
NExT-QA [18] Daily life M 44 5.4 V, A 52 Yes Yes Yes Yes MC & OE
KnowIT-VQA [19] TV shows M 60 12 V, A, S 24 Yes No Yes Yes MC
AGQA [20] Social Media A 30 9.6 V, A 192 Yes No Yes Yes OE
AVQA [21] Social Media M 60 9.3 V, A 57.3 Yes No Yes Yes MC
STAR [22] Daily life A - 22 V 60 Yes No No Yes MC
Env-QA [23] Egocentric Video M 20 23.3 V 85.1 Yes No No Yes MC
FIBER [24] Daily life M 10 28 V, A, S 2 No Yes No Yes OE
HumorQA Daily life M 7 1.8 V, A 141.3 Yes Yes Yes Yes OE & FT
CreativeQA Performance M 48 0.9 V, A 78.7 Yes Yes Yes Yes OE & FT
MagicQA Magic M 10 1.6 V, A 91.9 Yes Yes Yes Yes OE & FT
FunQA Surprising videos M 19 4.3 V, A 312 Yes Yes Yes Yes OE & FT
Video Question Answering Benchmarks While the visual question answering (VQA) task focuses on enhancing models’ ability in image comprehension [25, 26, 27], video question answering (VideoQA) shifts the attention towards video comprehension. VideoQA is generally more challenging than VQA as it requires a comprehensive understanding of visual content, utilization of temporal and spatial information, and exploration of relationships between recognized objects and activities [14]. To address the VideoQA task, the research community has introduced various benchmarks. As depicted in Table 1, Most commonly used VideoQA datasets are sourced from human-centric videos like movies [13], TV shows [15, 16, 19], and social media [12, 17, 18, 20, 21, 22, 24], and there are also object-centric datasets of game videos [11], synthetic videos [14] and egocentric videos [23]. MovieQA [13] and TVQA [15] are commonly employed by VideoQA methods, which put forward tasks related to temporal and causal reasoning. However, they rely heavily on dialogue comprehension and textual plot summaries, which severely limits the challenge of visual reasoning. TGIF-QA [12] uses animated GIFs to challenge spatial-temporal reasoning, but as most GIFs are short videos of 3 seconds, and its tasks mainly focus on action description, TGIF-QA lacks complex reasoning evaluation ability. When most datasets use multiple choice questions as QA tasks, some methods, such as NExT-QA [18], try to join open-ended questions. NExT-QA mainly focuses on daily life videos, but the open-ended answers are mostly simple sentences containing only a few words. To sum up, most existing methods focus on ordinary videos, lack of understanding of intriguing or unexpected videos, and advanced reasoning tasks such as generating complete explanatory texts of videos remain to be explored.
Video Question Answering Solutions Earlier studies have explored various models, including LSTMs and graph-based neural networks, to capture cross-modal information [28, 29]. With the advent of Transformers, video understanding models like ClipBERT [30] and CoMVT [31] emerged, focusing on the comprehension of specific frames within a video. Subsequent models, such as Violet [32], extended their ability to encompass temporal and spatial information. However, these methods have primarily been applied to short videos. In the realm of long videos, MIST [33] stands out by achieving state-of-the-art performance and excelling in terms of computation efficiency and interpretability. Furthermore, recent Vision Language Models (VLMs) such as [34, 35, 36] have showcased remarkable video understanding capabilities.
Counter-Intuitive Benchmarks While many current computer vision benchmarks primarily focus on understanding commonsense content, there is a growing interest in addressing the realm of counter-intuitiveness. Several emerging benchmarks and models cater to this domain, such as Whoops [37], which emphasizes weird, unusual, and uncanny images, and MemeGraphs [38], which revolves around memes featuring humor and sarcasm. Furthermore, some work even challenges models to comprehend complex multimodal humor in comics [39]. In the realm of large vision-language models, exemplified by GPT-4 [40], there is a particular focus on showcasing their ability to provide explanations for funny pictures. However, when it comes to videos, existing datasets exploring humor in TV shows or comedy tend to heavily rely on audio and narrative cues [8, 9, 10], with visual clues playing a comparatively lesser role.
3 FunQA Dataset
In this section, we provide a detailed explanation of the design principles that guided the creation of the FunQA dataset and its subsets. Additionally, we introduce our novel VideoQA tasks tailored for FunQA, and FunQA data statistics in Figure 2. We introduce our construction pipeline in the end.
(a) Word cloud of HumorQA
(b) Word cloud of CreativeQA
(c) Word cloud of MagicQA
(d) Number of videos and split
(e) Total video length
(f) Number of QA pairs
(g) Time Dist. of localization task
(h) Avg. word count of answers
(i) Anno. consensus evaluation
Figure 2: Statistics of FunQA Dataset. FunQA consists of three subsets, each corresponding to different video types, and is annotated with free-text QA pairs. The first row displays word clouds representing critical annotations for each subset. The second row provides key dataset statistics, including the number of videos for different splits, video length, and QA pair count for three subsets. In the last row, (g) highlights the high-frequency time span of the answer for localization questions in red, (h) shows the average word count of answers, and (i) presents the percentage of consensus between annotators for the same answer in a sampled set.
3.1 Video Selection
In constructing the dataset, we adhered to three principles to address the challenges in video understanding capabilities: our dataset, FunQA, is visual centered and emphasizes counter-intuitive reasoning, spatial-temporal reasoning. Based on these principles, we collect 4365 videos from 3 different art genres and created three subsets: HumorQA, CreativeQA, and MagicQA.
HumorQA HumorQA composed of 1,769 meticulously curated web videos, serves as a unique source of insight into human humor comprehension. Notably, it contains the shortest average video length of 7s among the three subsets. We believe that the human process of understanding humor is complex and deep, requiring a holistic understanding of the video and adding a degree of common sense to it. Psychological research has demonstrated that humor arises from the incongruity [41, 42] between reality and expectations, flourishing with the skillful juxtaposition and transformation of events [43, 44, 45]. This makes humorous videos a valuable asset for the VideoQA dataset, anticipated to enhance a model’s proficiency in integrating information and performing deep reasoning.
CreativeQA CreativeQA is a collection of 927 videos averaging 48s in length from a TV show called Kasou Taishou [46]. This program, showcasing original and novel skits performed by various amateur groups and judged by a panel, boasts a strong creative flair [47]. The essence of the show lies in using a mix of people and props to mimic reality, with audiences deriving pleasure from information integration and comparison. We anticipate that the imitation nature of the show will challenge the model’s capacity for information extraction, while the longer video length and need for understanding creativity will put to test the model’s comprehension of spatial-temporal information.
MagicQA MagicQA encapsulates 1672 magic performance videos sourced from across the web, spanning various genres like camera magic, close-up magic, and stage magic. The essence of magic revolves around the creation of seemingly impossible illusions [48], employing diverse effects such as disappearance, creation, and transformation. These illusions are infused with abundant spatial-temporal information. Through this dataset, we aim to empower the model to not only track the ensuing changes in objects but also unravel the underlying mechanics [49] of these transformations.
3.2 Task Definition
To comprehensively evaluate the model’s ability to understand surprising videos, we designed the following 4 types of tasks for each subset:
Counter-intuitive Timestamp Localization Task The localization task is the base task to assess the model’s comprehension abilities. It involves localizing counter-intuitive segments within the video, answers expressed in either seconds or frames. This task serves as the basis for the subsequent two main tasks in the three subsets, where the focus shifts to locating moments of humor, creativity, and magical effects, respectively. Successfully completing this task demands the model’s understanding of the video’s overall content, incorporating both temporal and spatial information.
Detailed Description Task The description task aims to evaluate the model’s information extraction capabilities, serving as a fundamental aspect of video understanding. Across all three subsets, this task requires providing a free-text answer that describes the selected moment. Furthermore, this task allows for analysis of how the model extracts information and generates answers for subsequent tasks. By examining the model’s performance in this task, we gain insights into its ability to extract relevant information and generate meaningful responses.
Counter-intuitiveness Reasoning Task The reasoning task is designed to test the model’s ability to reason about the video, and in the three subsets, this question is Why Humorous, Why creative, and Why counter-intuitive and the answer is a free-text explanation. This task is very difficult and involves the model’s deep reasoning ability; it requires the model to give a complete explanation using information from the entire video and its own common sense.
Higher Level Tasks In addition to the three main tasks, we design higher-level tasks to enhance the model’s inference abilities on counter-intuitive videos. Title Task in HumorQA and CreativeQA requires generating a concise title summarizing the video’s content. Creative Scoring Task in CreativeQA involves rating the creativity of videos between 0 and 20. Magic Method Task in MagicQA requires the model to explain clearly the rationale behind the magic, and its purpose is to test the model’s ability to reason more deeply. To ensure the accuracy of the answers, this task is only partially annotated and appears only in the test set, details of which can be found in Appendix A.1.
In addition, some visual QA benchmarks [37] have begun to adopt free-text answers; this flexible answer format enables multimodal LLMs to go beyond multiple choice options (Large Language Models) [35, 36]; for Video QA, however, free-text are less common. To bridge this gap, we also consider some free-text answer evaluations, demanding enhanced video comprehension capabilities from the model. It will also assess the model’s ability to generate reasonably lengthy textual responses.
3.3 Dataset Statistics
FunQA contains 4,365 counter-intuitive video clips and 311,950 question-answer pairs, the total length of these videos is 23.9h and the average length of video clips is 19 seconds. FunQA consists of three fine-grained subsets, each one containing well-designed tasks. The specific numbers and split of videos can be seen in Figure 2 (d). Each subset’s video lengths can be seen in Figure 2 (d). The specific number of QA pairs for each task can be seen in Figure 2 (e).
For our localization task, the timestamp heat map for the three different types of videos can be seen in Figure 2 (g), which shows the high-frequency time span of the answer. For the description and reasoning tasks, the average length of the words in their free-text answers reached 34.24, which is much longer than existing VideoQA datasets (e.g., 8.7 in Activity-QA [50], 11.6 in NExT-QA [18]). The specific word count of each task is shown in Figure 2 (h). FunQA has a well-established annotation process and high annotation quality, the result of our annotation consensus evaluation can be seen in Figure 2 (i). For each video category, more than 90% of the annotations exhibit a high level of consensus, with only 1% of the content showing low consensus. Approximately 8% of the data demonstrates variations in consensus, thus highlighting the objectivity of our dataset.
HumorQA, CreativeQA, and MagicQA word clouds are shown in Figure 2 (a-c). More statistics and FunQA full word cloud are given in Appendix A.3.
3.4 Dataset Construction Pipeline
FunQA dataset construction pipeline was in three stages: Pre-processing, Manual Annotation, and Post-Processing. The whole process took about 900 hours with over 50 highly educated undergraduates as part-time annotators, and we paid crowd-workers a target of ¥25/hr. More details of the dataset construction pipeline can be seen in Appendix A.1.
Pre-processing Initially, we crawled videos from TikTok, Bilibili, and YouTube (these videos are a collection of surprising videos). Then we performed a two-stage manual cleaning and cutting process on the collected videos to ensure counter-intuitive features and video quality and to exclude non-ethical content and sensitive information, resulting in a cut video clip.
Manual Annotation We annotated the videos according to the characteristics of different tasks design in Chinese. We screen and train the annotators to ensure the accuracy and high quality of the annotation, and finally produce the original annotated files. After the first round of annotation, we conducted a secondary annotation of 10% of the tasks and performed Consensus Evaluation to ensure the objectivity of our annotations.
Post-processing Based on our carefully designed tasks and high-quality annotations, we expanded our dataset using GPT-3.5. Firstly, we automatically translated the Chinese annotations into English. Subsequently, we generated more QA pairs that were faithful to the original ideas but presented differently. This not only made FunQA multilingual but also increased the number of QA pairs to 312K. Additionally, we created diverse task types, such as FunQA multiple-choice and FunQA dialogue. More details are given in Appendix A.2 and Appendix B.
4 Limitations and Future Work
This paper has two limitations. 1) Current FunQA dataset mainly includes video-level data and annotations, but denser annotations (e.g., like PVSG [68]) can be developed to explore more possibilities of video reasoning. Examples include detailed spatial and temporal annotations, such as captions corresponding with specific time axes and annotations of object level. 2) The raw annotations are completed by the annotator in Chinese. In the process of translating into English, we first use GPT to polish and supplement Chinese annotations, making the text as thorough as possible, as short or incomplete text may result in misunderstanding. However, there may still be differences due to cultural diversities between the two languages.
In the future, we expect to expand our dataset with denser and more diverse annotations. Also, new metrics will be explored to better evaluate models’ performance, especially in open-ended questions which lack in-depth metrics. Finally, we hope to provide directions for advancing models toward deeper video reasoning. Specific ideas include that Otter can sample more frames from the input video to improve causal reasoning capacity. Furthermore, in the future, when refining tuning datasets such as MIMIC-IT [69] for visual language models, it may be beneficial to incorporate the FunQA training set. However, to ensure fair comparisons and prevent data leakage, it is advised that future research does not utilize the FunQA testing set.
Acknowledgments and Disclosure of Funding
This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221- 0012), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
Appendix A More details of FunQA dataset
A.1 Dataset Construction Pipeline
Figure 3: Example of Magic Method Task. During the annotation process, we discovered that even as humans, it is difficult to fully understand the complete principles behind the implementation of magic tricks right from the beginning.
Figure 4: FunQA Video Selection and Clip Principle. We have a zero-tolerance policy regarding the inclusion of offensive content in our dataset. During the video sourcing process (video selection and video clipping stages), we ensure that such content is completely eliminated.
Figure 5: Examples of incorrect annotations and the suggested modifications. Here we clarify the meanings of different cell colors. Red: Errors made by the annotator during the annotation process. Yellow: Types of annotation errors. Green: Reasons for the errors and suggested corrections.
Figure 6: GPT prompt for Translation and augmentation. Under our carefully designed prompt, GPT-3.5 can automatically generate English translations and five synonymous sentences from our high-quality Chinese annotations, thereby expanding FunQA dataset.
Pre-processing & Qualification For videos related to humor and magic, we downloaded them from different streaming platforms, mostly in the form of compilations. For Creative videos, we downloaded 26 episodes publicly available from Tokyo TV in Japan. We provided rigorous training to the annotators to ensure high-quality video clips in the final compilation. Annotators who successfully completed the Clip task according to the requirements are considered qualified and can proceed to the next stage of annotation.
Training & Annotation We conducted systematic training for all annotators who passed the previous round of annotation, focusing on different tasks:
For the Counter-intuitive Timestamp Localization Task (H1, C1, and M1), the annotation format is a pair of numbers enclosed in square brackets, [xxxx, xxxx]. We asked the annotators to record the time intervals in which they felt pleasure (or amusement or shock) while watching the video.
For the Task Detailed Description Task (H2, C2, and M2), we requested objective descriptions of what happened at [xxxx, xxxx], emphasizing a "what you see is what you get" approach. It is important to note that the annotations should only cover the selected time intervals and should not include subjective adverbs (such as vividly, vividly, or wildly). When describing characters or objects, be concise and add modifiers if there is ambiguity. An example of a poor label is "a man and a woman," while a relatively better example is "a man wearing a red hat and a woman wearing an apron."
For the Counter-intuitiveness Reasoning Task (H3, C3, and M3), explain why the video is interesting in the context of the overall content. This part requires interpretive answers based on analysis, reasoning, and prior knowledge to explain why the video is counter-intuitive.
Specifically for the Magic Method Task (M4), we found that most annotators were not professional magicians, and even when watching instructional magic videos, it was challenging for them to provide complete and reasonable annotations for magic methods. Therefore, for M4, we only provided a small number of examples, which do not exist in the FunQA dataset. Figure 3 shows an example of M4.
We conducted strict quality reviews in real-time for the annotations, including but not limited to checking if the annotations meet the requirements, if the selected videos contain sensitive content, and if the annotation texts contain sensitive content. The annotation principles (video selection and video clip, etc) can be seen in Figure 4. Figure 5 also illustrates some errors encountered during the annotation process and the suggested modifications provided. After completing all the annotations, we conducted a consensus evaluation study to assess the objectivity of our annotations. We randomly selected 10% of the videos and asked annotators to provide a consistency score (high consensus, medium consensus, low consensus) between their first and second annotations, considering all the previous annotations shown to them.
Copyright & License We respect the copyright of each video. For content downloaded from streaming platforms, we obtained permission from the respective uploaders. For Kasou Taisho, we submitted a fair use request to the official website of Tokyo TV in Japan.
We respect the personal identity information of everyone appearing in our videos and always strive to eliminate offensive content. By conducting strict reviews of the annotators and real-time quality checks during each annotation process, we ensure the absence of offensive content.
A.2 Language Augmentation
For each QA pair, we first asked GPT-3.5 to translate the previous Chinese answers into English, and then let GPT-3.5 give five separate answers with the same meaning but different linguistic expressions. Afterward, we filtered out incorrect and incomplete generations. Figure 6 shows the prompt we gave to GPT-3.5.
However, as mentioned in the main text in the limitation, the translation of GPT-3.5 will be ambiguous when the original text uses Chinese harmonics and some special Chinese meanings, and this problem mostly occurs in the Title Task, we screened such problems by hand and modified them, as shown in Figure 7.
Figure 7: Issues in translation caused by Chinese-English cultural differences. In this example of translating and expanding a humorous video title, GPT-3.5 failed to understand the meaning of the original Chinese title, and we filtered out such data and made manual changes.
A.3 Data Statistics
We collected raw videos from multiple video platforms, including short videos, long videos, and video clips, and the detailed data can be seen in Table 2. The word cloud shaped by FunQA logo of the all annotation word is shown in Figure 8.
Table 2: Statistics of the FunQA raw data.
Raw Data Type Source Videos Avg.len(s) Total.len(h)
Humor Daily Life (Human) Bilibili 351 182 15.8
TikTok 1296 14 5.18
Nature (Animal) Bilibili 230 133 8.52
Creative Performance Kasou Taishou 26 6060 43.77
MagicQA Close up Magic Youtube 765 96 20.40
Camera Magic 334 152 14.17
FunQA - - 3002 129 107.87
Figure 8: Word cloud of FunQA free-text answer.
Appendix B FunQA Extension Datasets
Our main objective in designing these extended datasets is to leverage our high-quality annotated data and provide a rich and suitable data format for models. In addition, we also wanted to test the capability of GPT-3.5 and the quality of the dataset, and it turns out that GPT-3.5 can expand our data tens of times after providing high-quality free-text annotations, and our FunQA dataset is extremely scalable.
Figure 9: GPT prompt for FunQA Multi-choice. We give GPT-3.5 prompt as system, and then enter the description and explanation of the video, and GPT-3.5 will output the data in the form of multiple-choice questions according to the prompt.
Figure 10: GPT prompt for FunQA Dialog. We give GPT prompt as system, and then input the description and explanation of the video, and GPT-3.5 will output the data in the form of dialogue according to the prompt.
B.1 FunQA Multi-choice Dataset
FunQA Multi-choice Dataset is prepared to provide training and testing for arbitrary models, in this dataset our QA pairs are in the form of multiple choice, the answer is a word, phrase, or short sentence, and the type of questions are all descriptions. An example of the prompt given to GPT-3.5 and the data generated can be seen in Figure 9.
B.2 FunQA Dialog Dataset
Most of the current LLMs are in the form of dialogues. To cater to their data input, we produced the FunQA Dialog dataset, in which we used GPT-3.5 to convert QA pairs into recursive dialogues with added context. In addition to this, we also let GPT-3.5 freely generate future dialogue development, which greatly expands the amount of information. An example of the prompt given to GPT-3.5 and the data generated can be seen in Figure 10.
Appendix C More Details of Experiment
C.1 Significance of New GPT-4 Based Metric
GPT-4 prompt design For each of the three tasks, we designed three prompts for scoring. For Detailed Description Task (H2, C3, and M2), we designed the prompt in five areas: text length, text content variation, text detail variation, logical text description, and linguistic ability. For Counter-intuitiveness Reasoning Task (H3, C3, and M3), we designed the prompt in six areas: expressiveness of language, the logic of response, the common sense of response, understanding of counter-intuition, differences in text detail, and length of text. For Title Task (H4 and C4), we used the description, comprehension, and title of the manually annotated video as a reference to score the new title. Each Prompt can be seen in Figure 11, 12, and 13.
Figure 11: GPT-4 prompt for Title Evaluation. We give GPT prompt as system, and then input a description, explanation, and two titles, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
Figure 12: GPT-4 prompt for Description Evaluation. We give GPT-4 prompt as system, and then input two descriptions, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
Figure 13: GPT-4 prompt for Explanation Evaluation. We give GPT prompt as system, and then input two explanations, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
Comparison between GPT-4 and Traditional Metrics The principle of the traditional metrics is relatively simple. The traditional metric principles are as follows.
BLEU-4 BLEU, full name is Bilingual Evaluation Understudy, is a commonly used machine translation evaluation metric. It evaluates how good a machine translation result is by comparing how well it matches the N-gram of one or more human-translated reference results, which is a sequence of N consecutive words. BLEU-4, i.e., evaluates how well a combination of two words (i.e., a binary) matches. BLEU introduces a correction factor, Brevity Penalty (BP), to avoid this problem, which penalizes machine translation if the result is shorter than the reference translation. The BLEU score is the geometric mean of the individual N-gram accuracy multiplied by the shortness penalty. That is, the score of BLEU takes into account the precision and length of the translation result.
ROUGE-L ROUGE, known as Recall-Oriented Understudy for Gisting Evaluation, is a commonly used evaluation method for tasks such as automatic digesting and machine translation. ROUGE is mainly evaluated by comparing the overlap between the generated abstracts and the reference abstracts. Among them, ROUGE-L is an important variant of ROUGE, where L stands for Longest Common Subsequence (LCS), i.e., the longest common subsequence. Unlike the n-gram, the longest common subsequence does not require consecutive occurrences of items in the sequence.
CIDEr CIDEr, known as Consensus-based Image Description Evaluation, is an evaluation metric specifically designed for evaluating image description (Image Captioning) tasks. The main advantage of CIDEr is that it can capture more detailed information because it uses TF-IDF weights to emphasize n-grams that occur frequently in manual annotation but are not common in all image descriptions.
BLEURT BLEURT, full name Bilingual Evaluation Understudy with Representations from Transformers, is an evaluation method based on the transformer model, specifically for evaluating the output quality of machine translation and natural language generation tasks. Unlike traditional evaluation metrics such as BLEU and ROUGE, BLEURT does not directly compare the n-gram match between generated text and reference text but uses pre-trained language models (e.g., BERT) to understand the semantic information of text. the advantage of BLEURT is that it can capture the deep semantic information of text, and can address some problems that are difficult to be handled by traditional evaluation metrics (e.g. synonym substitution, utterance rearrangement) giving reasonable evaluation.
It can be seen that the principles of the above traditional NLG metric are relatively simple and lack the ability to understand and evaluate the text with complex logic. In contrast, GPT-4, as an LLM, has a high ability to understand the text and can understand the text first and evaluate the similarity through prompt design.
Instability of GPT-4 responses During the experimental process using GPT-4 as the metric, we found that in a few cases, GPT-4 would provide different evaluations for the same content. In such cases, we would perform multiple evaluations and take the average. Figure 14 shows an example of GPT-4 generating different responses to the same content.
Figure 14: GPT-4 generates different responses to the same content. User Input is the two sentences input to GPT-4, which is used to judge the similarity. The GPT-4 Output 1 and GPT-4 Output 2 below are the answers given by GPT-4 in two different conversations, and you can see that the two assessments are different.
C.2 On Failure of the Timestamp Localization Task
There is no model that has the ability to answer the Timestamp Localization Task. Figure 15 shows exactly what is answered for each model.
Figure 15: Model responses on Timestamp Localization Task. As can be seen, the model gives answers to the Timestamp Localization Task still focusing on describing the video content and still does not answer the specific time period when asked to type the number of frames and seconds.
C.3 Implementation Details
mPLUG mPLUG is a multi-modal system employing independent image and text encoders, a cross-modal network, and a text generation decoder, which is trained through prefix language modeling loss to generate captions from connected image and prefix sub-sequence representations.
GIT GIT is a system with an image encoder and a text decoder; it processes multiple video frames independently, adds learnable temporal embeddings before concatenation, uses a contrastively pre-trained model for image encoding, and employs a transformer module for text prediction. We used the 14M version and used two models, GIT_LARGE_VATEX and GIT_LARGE_MSRVTT, which were fine-tuned on the video captioning task for the VATEX and MSRVTT datasets, respectively.
VideoChat VideoChat, specifically the VideoChat-13B version, is an end-to-end system for video comprehension that combines pre-trained models. It utilizes QFormer to generate video embeddings and then employs LLAMA-13B for multimodal understanding and outputs video text descriptions with timestamps. In the experiment, we used VideoChat-13B with the hyperparameters: beam search number = 1, temperature = 1, video segments = 8, and token = 512.
Video-ChatGPT Video-ChatGPT is a vision-language model with a video encoder and LLM. It generates answers using video embeddings and benefits from a data-centric, human-assisted annotation framework for high-quality video instructional data. In the experiment, we used Video-ChatGPT-7B with its hyperparameter: temperature = 0.2, and token = 512.
Otter The Otter model employs the OpenFlamingo training paradigm, utilizing pre-trained encoders for language (LLaMA-7B) and vision (CLIP ViT-L/14). In the fine-tuning process, Otter prioritizes the Perceiver resampler module while keeping the encoders frozen. In the training stage, we finetuned Otter on Dense Caption and FunQA for a total of 3 epochs each. In the experiment, we used two versions of Otter with the same hyperparameters: beam search number = 3, size of no-repeat-ngram = 0.2, and token = 256.
All of our experiments took five to six days on a single NVIDIA A40 48GB GPU.
C.4 More Examples
Figure 16 shows the responses of different models on CreativeQA and MagicQA.
Figure 16: Model responses on CreativeQA and MagicQA. For the description of the Creative video example, only VideoChat gives the key point of the bicycle, but its description also has many errors and omissions, and the remaining two models do not identify the bicycle. In the explanation task, the responses of all three models fail to clearly explain the creativity of this imitation performance. For the Magic video example, all three models perform very poorly in description and explanation, basically only answering the phone and the straw, but lacking the description and explanation of the magic effect.