### v2

# https://github.com/Jingkang50/FunQA

Abstract
Surprising videos, e.g., funny clips, creative performances, or visual illusions, attract significant attention. Enjoyment of these videos is not simply a response to visual stimuli; rather, it hinges on the human capacity to understand (and appreciate) commonsense violations depicted in these videos. We introduce FunQA, a challenging video question answering (QA) dataset specifically designed to evaluate and enhance the depth of video reasoning based on counter-intuitive and fun videos. Unlike most video QA benchmarks which focus on less surprising contexts, e.g., cooking or instructional videos, FunQA covers three previously unexplored types of surprising videos: 1) HumorQA, 2) CreativeQA, and 3) MagicQA. For each subset, we establish rigorous QA tasks designed to assess the model’s capability in counter-intuitive timestamp localization, detailed video description, and reasoning around counter-intuitiveness. We also pose higher-level tasks, such as attributing a fitting and vivid title to the video, and scoring the video creativity. In total, the FunQA benchmark consists of 312K free-text QA pairs derived from 4.3K video clips, spanning a total of 24 video hours. Moreover, we propose FunMentor, an agent designed for Vision-Language Models (VLMs) that uses multi-turn dialogues to enhance models’ understanding of counter-intuitiveness. Extensive experiments with existing VLMs demonstrate the effectiveness of FunMentor and reveal significant performance gaps for the FunQA videos across spatial-temporal reasoning, visual-centered reasoning, and free-text generation.
Figure 1: Overview of FunQA. FunQA comprises three subsets of surprising videos: 1) HumorQA, 2) CreativeQA, and 3) MagicQA. Each subset is associated with three common tasks: 1) counter-intuitive timestamp localization, 2) detailed video description, and 3) reasoning around counter-intuitiveness (see H1-3, C1-3, and M1-3). Furthermore, we offer higher-level tasks tailored for each video type, such as attributing a fitting and vivid title for HumorQA and CreativeQA (see H4, C4), etc.
1 Introduction
The charm of surprising videos, being funny, creative, and filled with visual illusions, offers enjoyment and attracts engagement from viewers. This type of media elicits positive surprise111c.f., negative surprise, e.g., a surprising medical bill. [46], a captivating emotion that stems not merely from perceiving surface-level visual stimuli, but rather, the innate ability of humans to understand and find delight in unexpected and counter-intuitive moments [40]. However, despite significant advancements in today’s computer vision models, the question remains: can video models “understand” the humor/creativity in surprising videos? Consider the humorous video depicted in Fig. 1 (left) as an example. We observe a woman in black holding two pot lids and clapping them together. The remaining three individuals are responsible for avoiding the pot lids. The first two people successfully dodge, but the third girl, in a panic, fails to avoid any hits and gets struck three times. The embarrassed demeanor of the third girl along with her final frustrated reaction, elicits laughter [^2] The hostility/superiority theory of humor posits that humor can arise from claiming superiority over someone or something [17, 2]; but alternate (more optimistic) theories of humor exist, [1] offers a survey.. While humans effortlessly recognize this as an unusual (and potentially entertaining) event, the reasoning required to holistically understand the scene is complex: a model needs to recognize that this is not a video depicting harm but rather girls engaging in playful pranks together, and discern that the comedic element arises from the stark contrast between the third girl being hit by the pot lids every time and the first two girls skillfully avoiding them.
Table 1: Comparison between FunQA and other existing benchmarks. Compared to other datasets, FunQA revolves around the captivating realm of interesting and counter-intuitive videos. The tasks within FunQA are specifically designed to challenge the vision capabilities of models, requiring strong skills in producing an in-depth description, interpretation, and spatial-temporal reasoning. Here we clarify the abbreviation in the table. For annotation type: [Uncaptioned image] denotes Manual Annotation and [Uncaptioned image] for Automatic Annotation; Avg Len denotes video average length; # Clips means number of video clips; VC for visual-centric, Des. for Description, Exp. for Explanation, STR for Spatial-temporal Reasoning, MC means Multiple Choice QA, and OE shows Open Ended QA with Average Word Count per response.
Dataset Domain [Uncaptioned image] or [Uncaptioned image] Video Question Answer
Avg Len # Clips # QA VC Des. Exp. STR MC OE
TGIF-QA [20] Social Media [Uncaptioned image] 3s 72K 165K 2.1
MSRVTT-QA [60] Social Media [Uncaptioned image] 15s 10K 244K 1.0
ActivityNet-QA [68] Social Media [Uncaptioned image] 180s 6K 58K 1.9
NExT-QA [59] Daily life [Uncaptioned image] 44s 5K 52K 2.6
Social-IQ [69] Daily life [Uncaptioned image] 99s 1K 8K N/A
MovieQA [54] TV shows [Uncaptioned image] 203s 7K 6K N/A
TVQA+ [33] TV shows [Uncaptioned image] 8s 4K 30K N/A
SUTD-TrafficQA [61] Traffic [Uncaptioned image] 5s 10K 623K N/A
MarioQA [44] Games [Uncaptioned image] 5s 188K 188K 2.0
CLEVRER [67] Synthetic Videos [Uncaptioned image] 5s 20K 305K N/A
FunQA (Ours) Surprising Videos [Uncaptioned image] 19s 4K 312K 34.2
While there have been some efforts to enhance computer vision models’ performance in Video Question Answering (VideoQA), these works have primarily focused on the common, less surprising videos found in existing VideoQA datasets. Examples of commonly employed VideoQA datasets include YouCook2 [72] which contains video clips from 2K cooking videos, Howto100M [41] which consists of only instructional videos. While there exist video datasets that explore the humor in TV shows [5, 18] and include tasks such as predicting laughter tracks [49], these tasks often heavily rely on audio and narrative cues, with visual clues might playing a lesser role. Beyond datasets centered on factual queries, it is worth noting that NExT-QA targets the explanation of video content, which is widely employed for evaluating reasoning abilities. However, it was found in the experiment (see Section LABEL:sec:compare_nextqa) that VLMs such as GPT-4V(ision) already achieved an accuracy of 80% on NExTQA. This demonstrates that with the development of VLMs, the demand for datasets with deeper reasoning capabilities and presenting greater challenges is increasing.
To revitalize the visual reasoning field and further improve model capabilities to identify and understand visual commonsense violations in videos, we introduce FunQA, an extensive and carefully curated VideoQA dataset comprising 4.3K surprising videos and 312K manually annotated free-text QA pairs. Unlike some VideoQA datasets that feature open-ended questions but short answers (e.g., an average of 2.6 words per answer in NExT-QA [59]), FunQA’s responses average 34.2 words in length. This significantly increases the demand for advanced video comprehension capabilities in the model. Therefore, here we use free-text QA to distinguish from open-ended QA. Our dataset consists of three subsets: 1) HumorQA, 2) CreativeQA, and 3) MagicQA. Each subset covers different sources and contents, but the commonality lies in their surprising nature, e.g., the unexpected contrasts in humorous videos, the intriguing disguises in creative videos, and the seemingly impossible performances in magic videos. Our experiments suggest that these surprising videos require different types of reasoning than common videos, as existing VideoQA methods perform poorly on the corpus. With FunQA, we hope to provide a benchmark that covers the popular, important, and sophisticated genre of counter-intuitive/surprising videos.
In FunQA, we formulate three rigorous tasks to measure models’ understanding of surprise: 1) Counter-intuitive timestamp localization: a model must identify the specific time period within a video when an unexpected event takes place. 2) Detailed video description: a model must generate coherent and objective descriptions of the video content, evaluating models’ fundamental video understanding capabilities. 3) Counter-intuitiveness reasoning: a model must generate concrete explanations of why the video is surprising. These tasks progressively assess the model’s ability to perceive, articulate, and reason about the counter-intuitive elements present in surprising videos. We also propose additional tasks that pose higher-level challenges, such as assigning an appropriate and vivid title to the video.
In continuation of our efforts to enhance models’ comprehension of surprising content, we introduce FunMentor, a specialized agent designed to boost counter-intuitive reasoning in VLMs. Operating like a seasoned coach in a variety show, FunMentor engages in detailed, multi-turn dialogues, honing the models’ responses to accurately grasp the essence of both amusing and astonishing content. Beyond just evaluation, FunMentor actively steers VLMs with precise prompts, fostering fluent, logical, and persuasive responses. Experiments have demonstrated its effectiveness in augmenting VLMs’ ability to comprehend.
To summarize our contributions:
1) New VideoQA Dataset: We build a large-scale dataset FunQA, which complements the existing VideoQA dataset with intriguing videos.
2) Novel and Challenging Tasks: We design a number of novel tasks that allow the model to explore previously untouched problems, such as timestamp localization, and reasoning around counter-intuitiveness. These tasks push video reasoning beyond superficial descriptions, demanding deeper understanding.
3) Novel Method named FunMentor: We propose a novel agent that refines the model’s understanding of counter-intuitiveness through multi-turn dialogues with VLMs.
4) Comprehensive Evaluation: We have done an extensive and comprehensive evaluation of cutting-edge baselines, giving the field an insight and future research direction.
2 Related Work
Video Question Answering Benchmarks While the visual question answering (VQA) task focuses on enhancing models’ ability in image comprehension [15, 73, 28], video question answering (VideoQA) shifts the attention towards video comprehension. VideoQA is generally more challenging than VQA as it requires a comprehensive understanding of visual content, utilization of temporal and spatial information, and exploration of relationships between recognized objects and activities [67]. To address the VideoQA task, the research community has introduced various benchmarks. As depicted in Table 1 (the complete table are shown in Appendix 0.A), most commonly used VideoQA datasets are sourced from human-centric videos like movies [54], TV shows [32, 33, 14], and social media [20, 69, 59, 16, 64, 58, 6], and there are also object-centric datasets of game videos [44], synthetic videos [67] and egocentric videos [12]. MovieQA [54] and TVQA [32] are commonly employed by VideoQA methods, which put forward tasks related to temporal and causal reasoning. However, they rely heavily on dialogue comprehension and textual plot summaries, which severely limits the challenge of visual reasoning. TGIF-QA [20] uses animated GIFs to challenge spatial-temporal reasoning, but as most GIFs are short videos of 3 seconds, and its tasks mainly focus on action description, TGIF-QA lacks complex reasoning evaluation ability. When most datasets use multiple choice questions as QA tasks, some methods, such as NExT-QA [59], try to join open-ended questions. NExT-QA mainly focuses on daily life videos, but the open-ended answers are mostly simple sentences containing only a few words. To sum up, most existing methods focus on ordinary videos, lack of understanding of intriguing or unexpected videos, and advanced reasoning tasks such as generating complete explanatory texts of videos remain to be explored.
Video Question Answering Methods Earlier studies have explored various models, including LSTMs and graph-based neural networks, to capture cross-modal information [37, 71]. With the advent of Transformers, video understanding models, like ClipBERT [31] and CoMVT [53] emerged, focusing on understanding specific frames within a video. Subsequent models like Violet [11], extended their ability to encompass temporal and spatial information. However, these methods have primarily been applied to short videos. For long videos, MIST [13] stands out by achieving state-of-the-art (SOTA) performance and excelling in terms of computation efficiency and interpretability. Furthermore, recent VLMs [34, 36, 43] have showcased remarkable video understanding capabilities.
Counter-Intuitive Benchmarks While many current computer vision benchmarks primarily focus on understanding commonsense content, there is a growing interest in addressing the realm of counter-intuitiveness. Several emerging benchmarks and models cater to this domain, such as Whoops [3], which emphasizes weird, unusual, and uncanny images, OOPS [9], which centers on recognizing and predicting unintentional events, and MemeGraphs [26], which revolves around memes featuring humor and sarcasm. Furthermore, some works even challenges models to comprehend complex multimodal humor in comics [19]. In the realm of large language models, exemplified by GPT-4 [47], there is a particular focus on showcasing their ability to provide explanations for funny pictures. However, regarding videos, existing datasets exploring humor in TV shows or comedy tend to heavily rely on audio and narrative cues  [5, 18, 49], with visual clues playing a comparatively lesser role.
3 The FunQA Dataset
(a) Number of videos and splits
(b) Number of QA pairs
(c) Time Dist. of Localization
(d) Consensus Evaluation
Figure 2: Statistics of FunQA Dataset. Figure (a) and (b) showcase vital statistics, including the number of videos for different videos, splits, and QA pairs count for three subsets. Figure (c) highlights the high-frequency time span of the answer for localization questions in red. Figure (d) presents the percentage of consensus between annotators for the same QA pair. The consensus is categorized into three levels High for consistent understanding, Medium for partial agreement but with mutual acknowledgment, and Low for complete disagreement.
In this section, we provide a detailed explanation of the design principles that guided the creation of the FunQA dataset and its subsets. Additionally, we introduce our novel VideoQA tasks tailored for FunQA, and FunQA data statistics in Figure 2. Toward the end we present our Construction Pipeline and Quality Control, highlighting our efforts in maintaining data quality and objectivity.
3.1 Task Definition
To comprehensively evaluate the model’s ability to understand surprising videos, we designed the following 4 types of tasks for each subset:
Counter-intuitive Timestamp Localization Task The localization task is the base task to assess the model’s comprehension abilities. It involves localizing counter-intuitive segments within the video, answers expressed in either seconds or frames. This task serves as the basis for the subsequent two main tasks in the three subsets, where the focus shifts to locating moments of humor, creativity, and magical effects, respectively. Successfully completing this task demands the model’s understanding of the video’s overall content, incorporating both temporal and spatial information.
Detailed Description Task The description task aims to evaluate the model’s information extraction capabilities, serving as a fundamental aspect of video understanding. This task requires providing a free-text answer that describes the selected moment. Furthermore, this task allows for analysis of how the model extracts information and generates answers for subsequent tasks. By examining the model’s performance in this task, we gain insights into its ability to extract relevant information and generate meaningful responses.
Counter-intuitiveness Reasoning Task The reasoning task is designed to test the model’s ability to reason about the video, and in the three subsets, this question is Why Humorous, Why creative, and Why counter-intuitive and the answer is a free-text explanation. This task is very difficult and involves the model’s deep reasoning ability; it requires the model to give a complete explanation using information from the entire video and its own common sense.
Higher Level Tasks In addition to the three main tasks, we design higher-level tasks to enhance the model’s inference abilities on counter-intuitive videos. Title Task in HumorQA and CreativeQA requires generating a concise title summarizing the video’s content. Creative Scoring Task in CreativeQA involves rating the creativity of videos between 1 and 20 (provided officially). Magic Method Task in MagicQA requires the model to explain clearly the rationale behind the magic, and its purpose is to test the model’s ability to reason more deeply. To ensure the accuracy of the answers, this task is only partially annotated and appears only in the test set, details of which can be found in Appendix 0.A.1.
3.2 Dataset Statistics
FunQA contains 4,365 counter-intuitive video clips and 311,950 QA pairs, the total length of these videos is 23.9h and the average length of each clips is 19 seconds. FunQA consists three fine-grained subsets, each one containing well-designed tasks. The specific numbers of videos and splits can be seen in Fig. 2 (a). The specific number of QA pairs for each task can be seen in Fig. 2 (b). For our localization task, the heat map for the three different types of videos can be seen in Fig. 2 (c), which shows the high-frequency time span of the answer. For the description and reasoning tasks, the average length of the words in their free-text answers reached 34.24, which is much longer than existing VideoQA datasets (e.g., 2.6 in NExT-QA [59]). FunQA has a well-established annotation process and high annotation quality, the result of our annotation consensus evaluation are illustrated in Fig. 2 (d). Impressively, over 90% of the annotations demonstrate a high level of consensus, while only 1% exhibit low consensus. This clearly underscores the objectivity and reliability of the FunQA dataset.
3.3 Dataset Construction Pipeline
FunQA dataset construction pipeline was in three stages: Pre-processing, Manual Annotation, and Post-Processing. The whole process took about 900 hours with over 50 highly educated undergraduates as part-time annotators. See Appendix 0.A.1 for more details on the dataset construction.
Pre-Processing Initially, we crawled videos from YouTube. Then we performed a two-stage manual cleaning and cutting process on the collection to ensure counter-intuitive features and video quality and to exclude non-ethical and sensitive content, resulting in video clips.
Manual Annotation We annotated the videos according to the characteristics of different task designs in Chinese. We screen and train the annotators to ensure the accuracy and high quality of the annotation, and finally produce the original annotated files. After the first round, we conducted a secondary round of 10% of the tasks and performed Consensus Evaluation to ensure the objectivity.
Post-Processing Based on our carefully designed tasks and high-quality annotations, we expanded our dataset using GPT-3.5. Firstly, we automatically translated the Chinese annotations into English. Subsequently, we generated more QA pairs that were faithful to the original ideas but presented differently. This not only made FunQA multilingual but also expanded its QA pair count to 312K. Additionally, we created diverse sub-dataset, FunQA-MC (multi-choice QA) and FunQA-DIA (dialogue QA). In addition, to focus on exploring the ability to handle counter-intuitive reasoning, we released FunQA-MC-R (a multi-choice version specifically containing counter-intuitive reasoning questions). More details are given in Appendix 0.A.2 and Appendix 0.B.
Table 2: Consensus Evaluation Experiment. This table shows the results from a random 10% sample of QA pairs, cross-validated by annotators to assess agreement with existing annotations. ‘Low,’ ‘Medium,’ and ‘High’ indicate the strength of the consensus. Consensus HumorQA CreativeQA MagicQA # Total Low 1 1 2 4 (1%) Medium 9 6 17 32 (8%) High 199 111 194 504 (91%) Table 3: Can FunQA be Solved Solely Based on Images? The left two columns show the average number of questions answerable or not by humans using only 8 static, uniformly selected frames. Dataset # Can # Cannot Cannot Rate HumorQA 7.8 32.2 80% CreativeQA 6.1 33.9 85% MagicQA 2.6 37.4 94% FunQA 16.5 103.5 86%
3.4 Quality Control
Minimal Errors and High Objectivity in FunQA We assure that every annotation included in the final release of FunQA has been subjected to rigorous multi-person, multi-round review processes. Furthermore, We did manual consensus evaluation on released FunQA dataset, randomly sampling 10% of the data from all three sub-datasets (HumorQA, CreativeQA, and MagicQA). As shown in Table 3.3, we get the 91% high consensus.
FunQA Emphasis on Temporal Dynamics FunQA requires a strong emphases on temporal dynamics rather than solely on few frames of images. To prove that, we did the quantitative human experiments - We randomly selected 40 videos from each of the three sub-datasets, totaling 120 videos. For each video, we sampled 8 frames evenly. We enlisted 10 individuals who had not seen any FunQA videos before. We had them view the sequence of 8 consecutive frames and then watch the original video along with its annotations. They were asked to determine whether they can understand and answer the counter-intuitive understanding of the original video solely based on the images. Nearly 86% people thought that FunQA cannot be solved only by images, as shown in Table 3.3.
4 Limitations and Future Work
This paper has two limitations. 1) The current FunQA dataset primarily contains video-level data and annotations. There is potential for enhanced video reasoning through denser annotations, akin to PVSG [63], which might include detailed spatial, temporal, and object-level annotations. 2) The initial annotations were made in Chinese and later translated into English. While GPT was used to refine and complete the Chinese text, ensuring comprehensiveness and correctness, differences due to cultural differences between the languages might still persist.
Looking ahead, we plan to enrich FunQA with more detailed and varied annotations. We aim to develop new metrics for a more accurate evaluation of models, particularly for open-ended questions. Our goal is to steer models towards deeper video reasoning. However, to ensure fair comparisons and prevent data leakage, it is advised that future research does not utilize the FunQA testing set.
5 Acknowledgments and Disclosure of Funding
This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221- 0012), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
Appendix 0.A More Details of FunQA Dataset
0.A.1 Dataset Construction Pipeline
Video Selection In constructing the dataset, we adhered to three principles to address the challenges in video understanding capabilities: our dataset, FunQA, is visual centered and emphasizes counter-intuitive reasoning, spatial-temporal reasoning. Based on these principles, we collect 4365 videos from 3 different art genres and created three subsets: HumorQA, CreativeQA, and MagicQA.
HumorQA HumorQA composed of 1,769 meticulously curated web videos, serves as a unique source of insight into human humor comprehension. Notably, it contains the shortest average video length of 7s among the three subsets. We believe that the human process of understanding humor is complex and deep, requiring a holistic understanding of the video and adding a degree of common sense to it. Psychological research has demonstrated that humor arises from the incongruity [23, 42] between reality and expectations, flourishing with the skillful juxtaposition and transformation of events [30, 4, 25]. This makes humorous videos a valuable asset for the VideoQA dataset, anticipated to enhance a model’s proficiency in integrating information and performing deep reasoning.
CreativeQA CreativeQA is a collection of 927 videos averaging 48s in length from a TV show called Kasou Taishou [8]. This program, showcasing original and novel skits performed by various amateur groups and judged by a panel, boasts a strong creative flair [52]. The essence of the show lies in using a mix of people and props to mimic reality, with audiences deriving pleasure from information integration and comparison. We anticipate that the imitation nature of the show will challenge the model’s capacity for information extraction, while the longer video length and need for understanding creativity will put to test the model’s comprehension of spatial-temporal information.
MagicQA MagicQA encapsulates 1672 magic performance videos sourced from across the web, spanning various genres like camera magic, close-up magic, and stage magic. The essence of magic revolves around the creation of seemingly impossible illusions [45], employing diverse effects such as disappearance, creation, and transformation. These illusions are infused with abundant spatial-temporal information. Through this dataset, we aim to empower the model to not only track the ensuing changes in objects but also unravel the underlying mechanics [29] of these transformations.
Table A1: Comparison between FunQA and other existing benchmarks (Complete Version). Compared to other datasets, FunQA revolves around the captivating realm of interesting and counter-intuitive videos. The tasks within FunQA are specifically designed to challenge the vision capabilities of models, requiring strong skills in producing an in-depth description, interpretation, and spatial-temporal reasoning. Here we clarify the abbreviation in the table. For annotation type: [Uncaptioned image] denotes Manual Annotation and [Uncaptioned image] for Automatic Annotation; Avg Len denotes video average length; # Clips means number of video clips; VC for visual-centric, Des. for Description, Exp. for Explanation, STR for Spatial-temporal Reasoning, MC means Multiple Choice QA, and OE shows Open Ended QA with Average Word Count per response.
Dataset Domain [Uncaptioned image] or [Uncaptioned image] Video Question Answer
Avg Len # Clips # QA VC Des. Exp. STR MC OE
TGIF-QA [20] Social Media [Uncaptioned image] 3s 72K 165K 2.1
MSRVTT-QA [60] Social Media [Uncaptioned image] 15s 10K 244K 1.0
ActivityNet-QA [68] Social Media [Uncaptioned image] 180s 6K 58K 1.9
MSVD-QA [60] Social Media [Uncaptioned image] 10s 2K 51K 1.0
YouTube2Text-QA [66] Social Media [Uncaptioned image] 10s 10K 123K N/A
AGQA [16] Social Media [Uncaptioned image] 30K 10s 192K TBD
AVQA [64] Social Media [Uncaptioned image] 60s 9K 57K N/A
NExT-QA [59] Daily life [Uncaptioned image] 44s 5K 52K 2.6
Social-IQ [69] Daily life [Uncaptioned image] 99s 1K 8K N/A
STAR [58] Daily life [Uncaptioned image] - 23K 60K N/A
FIBER [6] Daily life [Uncaptioned image] 10s 28K 2K TBD
MovieQA [54] TV shows [Uncaptioned image] 203s 7K 6K N/A
TVQA [32] TV shows [Uncaptioned image] 76s 22K 153K N/A
TVQA+ [33] TV shows [Uncaptioned image] 8s 4K 30K N/A
KnowIT-VQA [14] TV shows [Uncaptioned image] 60s 12K 24K N/A
SUTD-TrafficQA [61] Traffic [Uncaptioned image] 5s 10K 623K N/A
MarioQA [44] Games [Uncaptioned image] 5s 188K 188K 2.0
CLEVRER [67] Synthetic Videos [Uncaptioned image] 5s 20K 305K N/A
Env-QA [12] Egocentric [Uncaptioned image] 20s 23K 85K N/A
HumorQA (Ours) Daily life [Uncaptioned image] 7s 2K 141K 28.2
CreativeQA (Ours) Performance [Uncaptioned image] 48s 1K 79K 59.1
MagicQA (Ours) Magic shows [Uncaptioned image] 10s 2K 92K 27.6
FunQA (Ours) Surprising Videos [Uncaptioned image] 19s 4K 312K 34.2
Figure A1: Example of Magic Method Task. During the annotation process, we discovered that even as humans, it is difficult to fully understand the complete principles behind the implementation of magic tricks right from the beginning.
Figure A2: FunQA Video Selection and Clip Principle. We have a zero-tolerance policy regarding the inclusion of offensive content in our dataset. During the video sourcing process (video selection and video clipping stages), we ensure that such content is completely eliminated.
Figure A3: Examples of incorrect annotations and the suggested modifications. Here we clarify the meanings of different cell colors. Red: Errors made by the annotator during the annotation process. Yellow: Types of annotation errors. Green: Reasons for the errors and suggested corrections.
Figure A4: GPT prompt for Translation and augmentation. Under our carefully designed prompt, GPT-3.5 can automatically generate English translations and five synonymous sentences from our high-quality Chinese annotations, thereby expanding FunQA dataset.
Figure A5: Issues in translation caused by Chinese-English cultural differences. In this example of translating and expanding a humorous video title, GPT-3.5 failed to understand the meaning of the original Chinese title, and we filtered out such data and made manual changes.
Pre-processing & Qualification For videos related to humor and magic, we downloaded them from different streaming platforms, mostly in the form of compilations. For Creative videos, we downloaded 26 episodes publicly available from Tokyo TV in Japan. We provided rigorous training to the annotators to ensure high-quality video clips in the final compilation. Annotators who successfully completed the Clip task according to the requirements are considered qualified and can proceed to the next stage of annotation.
Training & Annotation We conducted systematic training for all annotators who passed the previous round of annotation, focusing on different tasks:
For the Counter-intuitive Timestamp Localization Task (H1, C1, and M1), the annotation format is a pair of numbers enclosed in square brackets, [xxxx, xxxx]. We asked the annotators to record the time intervals in which they felt pleasure (or amusement or shock) while watching the video.
For the Task Detailed Description Task (H2, C2, and M2), we requested objective descriptions of what happened at [xxxx, xxxx], emphasizing a "what you see is what you get" approach. It is important to note that the annotations should only cover the selected time intervals and should not include subjective adverbs (such as vividly, vividly, or wildly). When describing characters or objects, be concise and add modifiers if there is ambiguity. An example of a poor label is "a man and a woman," while a relatively better example is "a man wearing a red hat and a woman wearing an apron."
For the Counter-intuitiveness Reasoning Task (H3, C3, and M3), explain why the video is interesting in the context of the overall content. This part requires interpretive answers based on analysis, reasoning, and prior knowledge to explain why the video is counter-intuitive.
Specifically for the Magic Method Task (M4), we found that most annotators were not professional magicians, and even when watching instructional magic videos, it was challenging for them to provide complete and reasonable annotations for magic methods. Therefore, for M4, we only provided a small number of examples, which do not exist in the FunQA dataset. Fig. A1 shows an example of M4.
We conducted strict quality reviews in real-time for the annotations, including but not limited to checking if the annotations meet the requirements, if the selected videos contain sensitive content, and if the annotation texts contain sensitive content. The annotation principles (video selection and video clip, etc) can be seen in Fig. A2. Fig. A3 also illustrates some errors encountered during the annotation process and the suggested modifications provided. We assured that every annotation included in the final release of FunQA has been subjected to rigorous multi-person, multi-round review processes. Specifically, every piece of released FunQA version annotated content underwent scrutiny by three distinct annotators. After completing all the annotations, we conducted a consensus evaluation study to assess the objectivity of our annotations. We randomly selected 10% of the videos and asked annotators to provide a consistency score (high consensus, medium consensus, low consensus) between their first and second annotations, considering all the previous annotations shown to them.
Copyright & License We respect the copyright of each video. Our data sharing approach is informed by the models established by previous studie like Kinetics [24], HD-VILA-100M [62], and others. Instead of providing the original raw data, we only supply the YouTube video IDs necessary for downloading the respective content.
We respect the personal identity information of everyone appearing in our videos and always strive to eliminate offensive content. By conducting strict reviews of the annotators and real-time quality checks during each annotation process, we ensure the absence of offensive content.
0.A.2 Language Augmentation
For each QA pair, we first asked GPT-3.5 to translate the previous Chinese answers into English, and then let GPT-3.5 give five separate answers with the same meaning but different linguistic expressions. Afterward, we filtered out incorrect and incomplete generations. Fig. A4 shows the prompt we gave to GPT-3.5.
However, as mentioned in the main text in the limitation, the translation of GPT-3.5 will be ambiguous when the original text uses Chinese harmonics and some special Chinese meanings, and this problem mostly occurs in the Title Task, we screened such problems by hand and modified them, as shown in Fig. A5.
0.A.3 Data Statistics
For a complete comparison between FunQA and other existing benchmarks, please see Table A1. We collected raw videos from multiple video platforms, including short videos, long videos, and video clips, and the detailed data can be seen in Table A2. The statics of FunQA dataset before GPT-3.5 extension can be seen in Table A3. The word cloud of the all annotation word is shown in Fig. 5(d) (a-d).
Table A2: Statistics of the FunQA raw data.
FunQA Subset Type Source # Videos Avg.len (s) Total.len (h)
Humor Daily Life (Human) Youtube 351 182 15.8
1296 14 5.18
Nature (Animal) 230 133 8.52
Creative Performance Youtube 26 6060 43.77
MagicQA Close up Magic Youtube 765 96 20.40
Camera Magic 334 152 14.17
FunQA - - 3002 129 107.87
Table A3: Statistics of the FunQA before GPT-3.5 extension.
Datasets Avg length (s) # Clips (K) # QA pairs (K) # QA per clip
HumorQA 7 1.8 7.2 4.0
CreativeQA 48 0.9 4.5 5.0
MagicQA 10 1.6 4.8 3.0
FunQA 19 4.3 16.5 3.8
(a) HumorQA.
(b) CreativeQA.
(c) MagicQA.
(d) FunQA.
Figure A6: Word cloud of FunQA free-text answer.
Appendix 0.B FunQA Extension Datasets
Our main objective in designing these extended datasets is to leverage our high-quality annotated data and provide a rich and suitable data format for models. In addition, we also wanted to test the capability of GPT-3.5 and the quality of the dataset, and it turns out that GPT-3.5 can expand our data tens of times after providing high-quality free-text annotations, and our FunQA dataset is extremely scalable.
Figure A7: GPT prompt for FunQA Multi-choice. We give GPT-3.5 prompt as system, and then enter the description and explanation of the video, and GPT-3.5 will output the data in the form of multiple-choice questions according to the prompt.
Figure A8: GPT prompt for FunQA Dialog. We give GPT prompt as system, and then input the description and explanation of the video, and GPT-3.5 will output the data in the form of dialogue according to the prompt.
0.B.1 FunQA-MC Dataset
FunQA-MC (Multi-choice) Dataset is prepared to provide training and testing for arbitrary models, in this dataset our QA pairs are in the form of multiple choice, the answer is a word, phrase, or short sentence, and the type of questions are all descriptions. An example of the prompt given to GPT-3.5 and the data generated can be seen in Fig. A7. For the FunQA-MC-R, which is a multi-choice version specifically containing counter-intuitive reasoning questions, we randomly chose 100 QA pairs from Task 2 ( Counter-intuitiveness Reasoning Task) to create the subset.
0.B.2 FunQA-DIA Dataset
Most of the current LLMs are in the form of dialogues. To cater to their data input, we produced the FunQA-DIA (Dialog) dataset, in which we used GPT-3.5 to convert QA pairs into recursive dialogues with added context. In addition to this, we also let GPT-3.5 freely generate future dialogue development, which greatly expands the amount of information. An example of the prompt given to GPT-3.5 and the data generated can be seen in Fig. A8.
Appendix 0.C More Details of FunMentor
0.C.1 Real Fact Question
During the Real Fact Collection process, FunMentor relies on designing questions that VLMs can answer accurately and objectively to gain knowledge of the video’s basic information. The specific questions are shown in Fig. A9.
Figure A9: Example of Real fact question by FunMentor.
Figure A10: Prompt Design of FunMentor in multi-round dialogue. On the left are the prompts from FunMentor, which incorporate the information from collected Real Facts and conduct Answer Judgement to generate suggestions.
0.C.2 Prompts Design
Based on the QA pairs collected as mentioned, FunMentor engages in multi-round dialogues with VLMs through our predefined prompts. This process involves providing increasingly precise instructions, guiding the VLM towards correctly answering the question. Specific examples of this are shown in Fig. A10.
Table A4: Main Results on FunQA Benchmark (Complete Version). H2, C2, M2 represent the detailed video description task, and H3, C3, M3 represent reasoning around counter-intuitiveness. For the higher-level tasks, H4, C4 involve attributing a fitting and vivid title. The responses for all these tasks in free-text format. We use the following metrics: BLEU-4 / ROUGE-L / CIDEr (shown in the first row) and BLEURT / GPT-4 (shown in the second row) for evaluation. C5 represents scoring the video creativity, and the metric is the Accuracy between the predicted score and the official score. We tested the caption-based and instruction-based models. Here we clarify the abbreviation in the table. L.M.: GIT_LARGE_MSRVTT; L.V.: GIT_LARGE_VATEX; D.C. means finetuned on Dense Caption; FunQA means finetuned on FunQA.
HumorQA CreativeQA MagicQA
Task H2-Des. H3-Rea. H4-Title C2-Des. C3-Rea. C4-Title C5-Score M2-Des. M3-Rea.
- Caption-based Model
mPLUG [35] 1.5 / 16.4 / 1.0 19.9 / 3.9 1.1 / 12.5 / 0.4 25.7 / 6.0 0.6 / 7.5 / 0.1 22.1 / 11.2 0.4 / 13.4 / 0.0 14.9 / 3.0 0.7 / 12.6 / 0.1 24.2 / 6.9 0.3 / 3.2 / 0.0 20.8 / 18.8 0.0 1.2 / 15.8 / 0.5 19.7 / 4.0 0.9 / 8.9 / 0.4 21.2 / 8.1
GIT (L.M.) [57] 0.5 / 12.8 / 0.2 22.4 / 3.6 0.0 / 0.0 / 0.0 0.0 / 0.0 1.1 / 7.7 / 0.7 17.0 / 8.9 0.0 / 6.40 / 0.0 14.4 / 3.8 0.0 / 0.0 / 0.0 0.0 / 0.0 0.3 / 1.5 / 0.2 7.1 / 11.3 0.0 0.2 / 11.2 / 0.1 19.4 / 8.2 0.0 / 0.0 / 0.0 0.0 / 0.0
GIT (L.V.) [57] 1.2 / 16.9 / 0.6 33.3 / 4.0 0.0 / 0.0 / 0.0 0.0 / 0.0 1.0 / 8.8 / 0.7 25.9 / 10.0 0.1 / 8.30 / 0.0 20.5 / 4.2 0.0 / 0.0 / 0.0 0.0 / 0.0 0.5 / 2.8 / 0.4 10.5 / 12.0 0.0 0.6 / 13.7 / 0.1 29.8 / 8.6 0.0 / 0.0 / 0.0 0.0 / 0.0
- Instruction-based Model
VideoChat [36] 0.5 / 13.7 / 0.0 44.0 / 17.9 0.5 / 13.5 / 0.0 45.4 / 31.9 0.8 / 5.1 / 0.5 20.2 / 31.7 0.3 / 7.50 / 0.0 21.7 / 5.9 0.3 / 7.70 / 0.0 22.8 / 17.7 0.2 / 1.2 / 0.2 7.3 / 31.1 67.5 0.6 / 15.5 / 0.0 47.4 / 8.2 0.3 / 9.2 / 0.0 43.1 / 44.6
Video-ChatGPT [43] 0.5 / 14.0 / 0.1 39.9 / 24.3 0.7 / 12.4 / 0.1 40.1 / 24.9 0.4 / 3.2 / 0.2 36.5 / 41.2 1.1 / 19.8 / 0.2 45.8 / 6.6 0.8 / 17.3 / 0.1 45.2 / 9.1 0.2 / 1.9 / 0.2 30.9 / 48.8 85.4 0.7 / 20.8 / 0.0 50.8 / 11.2 0.5 / 11.3 / 0.0 43.3 / 40.4
mPLUG-Owl [65] 0.5 / 13.3 / 0.2 44.5 / 10,7 0.5 / 13.2 / 0.2 47.3 / 35.0 0.6 / 5.3 / 0.0 29.8 / 48.8 0.8 / 16.5 / 0.1 43.0 / 5.0 0.7 / 13.3 / 0.0 44.7 / 10.6 0.1 / 1.1 / 0.1 23.9 / 36.3 66.7 0.4 / 16.7 / 0.0 46.4 / 8.6 0.5 / 8.4 / 0.0 43.9 / 30.9
Video-LLaMA [70] 0.3 / 12.1 / 0.1 48.4 / 7.7 0.5 / 12.6 / 0.0 42.9 / 29.0 0.5 / 0.4 / 0.0 46.5 / 34.1 0.7 / 8.80 / 0.1 45.5 / 7.2 0.4 / 9.70 / 0.0 41.1 / 17.2 0.3 / 0.8 / 0.1 42.3 / 31.2 64.2 0.4 / 14.3 / 0.0 50.1 / 10.2 0.5 / 7.9 / 0.0 39.0 / 28.0
Otter (D.C.) [34] 1.1 / 14.3 / 0.4 30.2 / 7.7 1.2 / 14.2 / 0.4 32.3 / 28.3 0.5 / 5.4 / 0.1 21.7 / 20.0 0.5 / 13.8 / 0.1 28.7 / 1.7 1.0 / 16.8 / 0.2 32.9 / 7.9 0.3 / 2.3 / 0.1 17.7 / 36.3 45.0 1.0 / 15.0 / 0.3 32.5 / 2.1 1.1 / 12.8 / 0.2 27.3 / 36.8
Otter (FunQA) [34] 1.5 / 18.1 / 0.9 38.4 / 8.9 1.3 / 15.4 / 0.5 42.6 / 31.7 0.8 / 5.9 / 0.5 47.5 / 32.1 1.5 / 19.6 / 0.5 40.0 / 7.3 2.2 / 21.2 / 0.5 41.1 / 8.8 0.3 / 4.3 / 0.3 44.5 / 38.8 69.4 2.6 / 23.8 / 1.6 44.7 / 10.3 3.4 / 20.3 / 2.6 44.5 / 47.5
Video-ChatGPT [43] + FunMentor (Ours) 0.6 / 14.2 / 0.6 65.2 / 33.2 1.1 / 14.4 / 0.3 57.5 / 36.5 1.0 / 5.7 / 0.6 50.2 / 65.1 1.4 / 20.2 / 0.3 66.3 / 14.2 1.4 / 18.8 / 0.2 58.7 / 23.4 0.3 / 2.2 / 0.3 45.3 / 52.2 85.4 1.1 / 23.2 / 0.4 55.1 / 13.3 2.1 / 14.4 / 0.9 46.3 / 54.8
Otter (FunQA) [34] + FunMentor (Ours) 1.9 / 20.2 / 1.1 33.4 / 13.4 1.7 / 19.3 / 0.9 37.8 / 45.8 0.9 / 6.8 / 0.9 58.3 / 34.2 2.4 / 23.1 / 1.1 60.4 / 11.0 3.3 / 26.5 / 0.7 44.4 / 9.3 0.4 / 4.7 / 0.3 53.9 / 43.5 69.4 3.1 / 25.5 / 2.5 43.5 / 12.81 5.6 / 21.2 / 3.4 38.91 / 56.4
Appendix 0.D More Details of Experiment
The complete result of FunQA Benchmark (with traditional metircs’ scores) are shown in Table A4.
0.D.1 Significance of New GPT-4 Based Metric
GPT-4 prompt design For each of the three tasks, we designed three prompts for scoring. For Detailed Description Task (H2, C3, and M2), we designed the prompt in five areas: text length, text content variation, text detail variation, logical text description, and linguistic ability. For Counter-intuitiveness Reasoning Task (H3, C3, and M3), we designed the prompt in six areas: expressiveness of language, the logic of response, the common sense of response, understanding of counter-intuition, differences in text detail, and length of text. For Title Task (H4 and C4), we used the description, comprehension, and title of the manually annotated video as a reference to score the new title. Each Prompt can be seen in Fig. A11, A12, and A13.
Figure A11: GPT-4 prompt for Title Evaluation. We give GPT prompt as system, and then input a description, explanation, and two titles, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
Figure A12: GPT-4 prompt for Description Evaluation. We give GPT-4 prompt as system, and then input two descriptions, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
Figure A13: GPT-4 prompt for Explanation Evaluation. We give GPT prompt as system, and then input two explanations, the first one is our annotation, and the second one is the output of the model, and GPT-4 will evaluate the similarity between these two texts according to the prompt’s requirements.
Comparison between GPT-4 and Traditional Metrics The principle of the traditional metrics is relatively simple. The traditional metric principles are as follows.
BLEU-4 BLEU, full name is Bilingual Evaluation Understudy, is a commonly used machine translation evaluation metric. It evaluates how good a machine translation result is by comparing how well it matches the N-gram of one or more human-translated reference results, which is a sequence of N consecutive words. BLEU-4, i.e., evaluates how well a combination of two words (i.e., a binary) matches. BLEU introduces a correction factor, Brevity Penalty (BP), to avoid this problem, which penalizes machine translation if the result is shorter than the reference translation. The BLEU score is the geometric mean of the individual N-gram accuracy multiplied by the shortness penalty. That is, the score of BLEU takes into account the precision and length of the translation result.
ROUGE-L ROUGE, known as Recall-Oriented Understudy for Gisting Evaluation, is a commonly used evaluation method for tasks such as automatic digesting and machine translation. ROUGE is mainly evaluated by comparing the overlap between the generated abstracts and the reference abstracts. Among them, ROUGE-L is an important variant of ROUGE, where L stands for Longest Common Subsequence (LCS), i.e., the longest common subsequence. Unlike the n-gram, the longest common subsequence does not require consecutive occurrences of items in the sequence.
CIDEr CIDEr, known as Consensus-based Image Description Evaluation, is an evaluation metric specifically designed for evaluating image description (Image Captioning) tasks. The main advantage of CIDEr is that it can capture more detailed information because it uses TF-IDF weights to emphasize n-grams that occur frequently in manual annotation but are not common in all image descriptions.
BLEURT BLEURT, full name Bilingual Evaluation Understudy with Representations from Transformers, is an evaluation method based on the transformer model, specifically for evaluating the output quality of machine translation and natural language generation tasks. Unlike traditional evaluation metrics such as BLEU and ROUGE, BLEURT does not directly compare the n-gram match between generated text and reference text but uses pre-trained language models (e.g., BERT) to understand the semantic information of text. the advantage of BLEURT is that it can capture the deep semantic information of text, and can address some problems that are difficult to be handled by traditional evaluation metrics (e.g. synonym substitution, utterance rearrangement) giving reasonable evaluation.
WUPS The Wu-Palmer similarity (WPUS) measure calculates how similar two word senses are. It considers the depths of the two synsets in the WordNet taxonomies, along with the depth of the Least Common Subsumer (LCS). The WUPS metric has certain limitations that make it difficult to use in VQA tasks. First, certain words are very similar in vocabulary, but their meanings can be very different. The problem may arise with color. For example, if the answer to a certain question is white and the system predicts the answer to be black, that answer will still get a WUPS score of 0.92, which seems high. Another limitation is that WUPS cannot be used for answers to phrases or sentences because it always deals with rigid semantic concepts, which are most likely to be single words. More examples as shown in A14.
It can be seen that the principles of the above traditional NLG metric are relatively simple and lack the ability to understand and evaluate the text with complex logic. In contrast, GPT-4, as an LLM, has a high ability to understand the text and can understand the text first and evaluate the similarity through prompt design.
Figure A14: The responses of VLMs to NExT-OE are scored using the WUPS and GPT-4 based metric. It can be observed that Otter’s responses, whether correct or close to the answer, receive scores near zero under the WUPS metric. In contrast, HGA’s completely incorrect responses (e.g., answering “cat” when the correct answer is “dog”) score highly under WUPS. On the other hand, the scoring by GPT-4 appears to be accurate.
Instability of GPT-4 responses During the experimental process using GPT-4 as the metric, we found that in a few cases, GPT-4 would provide different evaluations for the same content. In such cases, we would perform multiple evaluations and take the average. Fig. A15 shows an example of GPT-4 generating different responses to the same content.
Figure A15: GPT-4 generates different responses to the same content. User Input is the two sentences input to GPT-4, which is used to judge the similarity. The GPT-4 Output 1 and GPT-4 Output 2 below are the answers given by GPT-4 in two different conversations, and you can see that the two assessments are different.
0.D.2 On Failure of the Timestamp Localization Task
There is no model that has the ability to answer the Timestamp Localization Task. Fig. A16 shows exactly what is answered for each model.
Figure A16: Model responses on Timestamp Localization Task. As can be seen, the model gives answers to the Timestamp Localization Task still focusing on describing the video content and still does not answer the specific time period when asked to type the number of frames and seconds.
0.D.3 Implementation Details
mPLUG mPLUG is a multi-modal system employing independent image and text encoders, a cross-modal network, and a text generation decoder, which is trained through prefix language modeling loss to generate captions from connected image and prefix sub-sequence representations.
GIT GIT is a system with an image encoder and a text decoder; it processes multiple video frames independently, adds learnable temporal embeddings before concatenation, uses a contrastively pre-trained model for image encoding, and employs a transformer module for text prediction. We used the 14M version and used two models, GIT_LARGE_VATEX and GIT_LARGE_MSRVTT, which were fine-tuned on the video captioning task for the VATEX and MSRVTT datasets, respectively.
VideoChat VideoChat, specifically the VideoChat-13B version, is an end-to-end system for video comprehension that combines pre-trained models. It utilizes QFormer to generate video embeddings and then employs LLAMA-13B for multimodal understanding and outputs video text descriptions with timestamps. In the experiment, we used VideoChat-13B with the hyperparameters: beam search number = 1, temperature = 1, video segments = 8, and token = 512.
Video-ChatGPT Video-ChatGPT is a vision-language model with a video encoder and LLM. It generates answers using video embeddings and benefits from a data-centric, human-assisted annotation framework for high-quality video instructional data. In the experiment, we used Video-ChatGPT-7B with its hyperparameter: temperature = 0.2, and token = 512.
Otter The Otter model employs the OpenFlamingo training paradigm, utilizing pre-trained encoders for language (LLaMA-7B) and vision (CLIP ViT-L/14). In the fine-tuning process, Otter prioritizes the Perceiver resampler module while keeping the encoders frozen. In the training stage, we finetuned Otter on Dense Caption and FunQA for a total of 3 epochs each. In the experiment, we used two versions of Otter with the same hyperparameters: beam search number = 3, size of no-repeat-ngram = 0.2, and token = 256.
mPLUG-Owl mPLUG-Owl is a novel training paradigm that equips LLMs with multi-modal abilities through modularized learning of foundation LLM, a visual knowledge module, and a visual abstractor module. The training paradigm involves a two-stage method for aligning image and text, which learns visual knowledge with the assistance of LLM, while maintaining and even improving the generation abilities of LLM. In the second stage, languageonly and multi-modal supervised datasets are used to jointly fine-tune a low-rank adaption (LoRA) module on LLM and the abstractor module by freezing the visual knowledge module.
Video-LLaMA Video-LLaMA is built on top of BLIP-2 and MiniGPT-4. It is composed of two core components: (1) Vision-Language (VL) Branch (Visual encoder: ViT-G/14 + BLIP-2 Q-Former) and (2) Audio-Language (AL) Branch (Audio encoder: ImageBind-Huge). In VL Branch, a two-layer video Q-Former and a frame embedding layer (applied to the embeddings of each frame) are introduced to compute video representations. And In AL Branch, a two-layer audio Q-Former and an audio segment embedding layer (applied to the embedding of each audio segment) are introduced to compute audio representations.
0.D.4 Ablation Experiment about Audio
Audio is undoubtedly an important part of video, but existing models utilize audio in a very limited way, some of them don’t use audio as an input, and then get high scores on various lists. For example, Otter, mPLUG, video-ChatGPT and GIT in the paper, the outputs of these models are not affected by the audio, and only the visual information is understood; there are also some models that convert the audio into captions for text encoding and then input them into the model, such as VideoChat in the paper. However, The FunQA dataset is designed to emphasize the understanding of ‘fun’ through the visual aspect of videos. During the data collection phase, we focused on the visual elements, and in our experimental setup, we retained the audio of the videos when feeding them to the models. To investigate the role of audio, we conducted a set of ablation experiments by muting the audio during both training and testing phases. The results are shown in Table A5. This demonstrates that FunQA is an visual-centric dataset and audio could not ease the challenge of the benchmark.
Table A5: The ablation experiment about audio. The scores of the results after inputting the muted video, the traditional evaluation scores remain basically unchanged, and the scores of GPT-4 fluctuate very little (this is mainly due to the issues of instability of the evaluation method of GPT-4.
HumorQA CreativeQA MagicQA
H1
H2 H3 H4
C1
C2 C3 C4 C5
M1
M2 M3
VideoChat [36]
0.5 / 13.7 / 0.0 44.0 / 37.9 0.5 / 13.5 / 0.0 45.4 / 31.9 0.8 / 5.1 / 0.5 20.2 / 61.7
0.3 / 7.5 / 0.0 21.7 / 10.9 0.3 / 7.7 / 0.0 22.8 / 27.7 0.2 / 1.2 / 0.2 7.3 / 51.1 67.5
0.6 / 15.5 / 0.0 47.4 / 14.2 0.3 / 9.2 / 0.0 43.1 / 24.6
VideoChat (MUTED) [36]
0.5 / 13.5 / 0.0 40.0 / 34.5 0.5 / 12.5 / 0.0 43.7 / 31.9 0.8 / 5.1 / 0.5 22.4 / 57.8
0.3 / 7.4 / 0.0 24.0 / 10.0 0.4 / 7.7 / 0.0 20.2 / 30.5 0.2 / 1.2 / 0.2 6.8 / 45.3 67.5
0.6 / 15.0 / 0.0 48.0 / 12.2 0.3 / 9.2 / 0.0 44.0 / 23.4
Table A6: Human Performance on FunQA. 100 QA pairs (H2, H3, H4) from HumorQA in FunQA were randomly selected and answered by five different people. We use the following metrics: BLEU-4 / ROUGE-L / CIDEr / BLEURT / GPT-4 for evaluation.
Task H2 H3 H4
Human 1 6.0 / 28.8 / 6.5 / 46.6 / 82.4 2.2 / 20.8 / 1.1 / 45.0 / 85.2 0.8 / 9.2 / 1.1 / 20.5 / 76.6
Human 2 13.8 / 42.5 / 11.1 / 55.4 / 90.5 13.2 / 34.1 / 10.0 / 54.4 / 87.5 1.1 / 12.8 / 2.4 / 33.5 / 80.6
Human 3 0.7 / 16.5 / 0.5 / 36.2 / 77.5 0.7 / 14.7 / 0.2 / 44.1 / 69.0 0.2 / 3.1 / 0.3 / 15.4 / 76.8
Human 4 26.4 / 50.5 / 27.7 / 75.8 / 80.3 10.3 / 36.8 / 11.5 / 74.3 / 90.0 2.8 / 21.7 / 6.0 / 43.4 / 73.4
Human 5 3.5 / 16.7 / 3.7 / 48.0 / 86.6 2.4 / 16.4 / 1.8 / 47.9 / 90.3 0.8 / 10.5 / 0.8 / 23.9 / 88.2
Model (SOTA) 1.5 / 18.1 / 1.0 / 44.0 / 37.9 1.3 / 15.4 / 0.5 / 25.9 / 61.7 1.1 / 8.8 / 0.7 / 25.9 / 61.7
Human (Avg) 10.1 / 31.0 / 9.9 / 60.0 / 83.6 5.8 / 24.6 / 5.0 / 53.1 / 84.4 1.1 / 11.46 / 2.1 / 27.3 / 79.1
0.D.5 More Examples
We conducted experiments to compare machine versus human performance on FunQA. We randomly selected 100 HumorQA pairs from HumorQA and had five individuals who had not watched any FunQA videos provide answers to the questions (H2, H3, H4) for each video. Subsequently, we calculated the average score for each individual under the FunQA benchmark and the overall average score of human responses. The FunQA benchmark score is the selected SOTA score between models.The results are in Table A6.
0.D.6 More Examples
Fig. A17 shows the responses of different models on CreativeQA and MagicQA.
Figure A17: Model responses on CreativeQA and MagicQA. For the description of the Creative video example, only VideoChat gives the key point of the bicycle, but its description also has many errors and omissions, and the remaining two models do not identify the bicycle. In the explanation task, the responses of all three models fail to clearly explain the creativity of this imitation performance. For the Magic video example, all three models perform very poorly in description and explanation, basically only answering the phone and the straw, but lacking the description and explanation of the magic effect.
Appendix 0.E More Discussions
0.E.1 (Potential) More Essential Factors
Accurate understanding of the videos Through our analysis of failure cases, we’ve observed that many models struggle with accurately describing videos. While they might be adept at detecting objects within the videos, they often falter in comprehending the contextual relationship between sequential events. Such misinterpretations indicate that there’s a need for further exploration in this domain. The videos we’ve used can indeed serve as an invaluable dataset for probing video descriptions in depth.
Logic Reasoning The primary nature of our videos encompasses content that is counterintuitive and contradicts common sense. For models to understand these, it’s imperative they grasp the concept of "common sense." They must deduce what would typically transpire under normal circumstances and then use that perspective to humorously interpret the video. This necessitates the model to possess strong reasoning capabilities, especially when it comes to common sense reasoning.
Extra Knowledge - Sense of Humor To decipher the humor in a video, it’s plausible that understanding the fundamental principles of humor is crucial. This type of knowledge, along with many other tidbits of common sense and additional information, might enhance the model’s performance. Determining how to integrate valuable knowledge and discerning what counts as "valuable" are topics that warrant further exploration.
0.E.2 Potential Solutions
Model Size Increasing the number of parameters is a natural method to enhance the model’s performance. However, this approach comes with its own set of engineering challenges, requiring improvements in model optimization and deployment. We’re also curious about the relationship between the number of parameters and the performance on FunQA. This is an intriguing research point in itself, and our dataset can serve as an excellent test bed to further this exploration.
Data Quality We believe the emphasis for this task should be on data collection. Current trends with large and dynamic models have shown that having vast amounts of low-quality data isn’t as effective as a smaller quantity of high-quality data. Thus, we hope the community can discover the type of data that genuinely assists in understanding counterintuitive videos. This is a crucial research direction.
Training Strategy Studying training strategies is also essential. For instance, determining which type of data to start learning from, and understanding the significance of curriculum learning, among others.
Model Collaboration Ultimately, we might not need to solely focus on a single model to solve this problem. Perhaps multiple models collaboratively working on examples in an elegant manner could be a method to enhance performance. However, this approach might necessitate paying more attention to the overall efficiency of model implementation
0.E.3 The Emphasis of Temporal Dynamics in FunQA
HumorQA Example: Humor_Example.mp4
H2: An individual slipped on the staircase filled with ice and tumbled down to the very bottom, followed by a second person who also fell after witnessing the first person’s fall.
H3: The first person falling down the stairs step by step was already very funny, and the second person repeating the same mistake and falling down made it even more hilarious.
Explanation: As answered in H2, the main element of humor in this video is two people slipping down an icy staircase in a sequential order, and if the model does not make sense of the temporal information, it will not be able to give the sequential logic of the two people slipping down one after the other. It is also clear from the answer to H3 that the sequence of slipping backwards and forwards is one of the sources of humor, as the person at the back does not learn from the lesson of the former and slips down in a similar way, and this repetition of the wrongdoing brings humor to the situation.
CreativeQA Example: Creative_Exapmle.mp4
C2: At the center of the stage is a blue rectangular box. Following the little girl’s watering of the box, four individuals lying inside gradually lift their hands and legs, each at different paces and heights, until they stop and reveal green painted leaf-shaped objects on their limbs.
C3: The main creative element of this video is when the four individuals in a box, with varying heights and movement speeds, gradually raise their hands and feet to mimic the growing process of the carrot seeds planted by the little girl. The green leaf-shaped objects, tied and opened in advance with their hands and feet, are used to simulate the true sprouting of carrot sprouts, resulting in a lively and imaginative scene.
Explanation: The performance in this video mimics growing bean sprouts, the process of growing beans into sprouts by constantly watering the soil. Trying to describe the content of this video idea and explain where the idea came from requires an understanding of the temporal information. Analyzing the visual information together, the model has to understand the sequence of watering, bean sprouts growing from the soil, and bean sprouts growing taller and greener in order to answer the question accurately. This difference in understanding can be demonstrated on the C3 task, where without analyzing the temporal information, the optimal answer would be bean sprouts growing from the soil, and cause and effect logical relationships such as watering to grow bean sprouts would be ignored.